A planet of mass m is in an elliptical orbit about the sun with an orbital period T if A be the area of orbit then its angular momentum would be

Grade:12

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
Kepler's law of Areas states that a line that connects a planet to the sun sweeps out equal areas in equal times.
mathematically,
dA/dt =L/2m (as areal velocity=L/2m = dA/dt]
or L =2mA/T [Ans]

Regards,
Thank you,
Soumya Ranjan Mohanty.